Two cars move at a constant velocity in a straight line. If they move in opposite directions the cars will be 600m apart after 30s. If they move in the same direction they will be 25m apart after 25s. Calculate thier velocities in m/s.
Answer by ankor@dixie-net.com(22740)

Two cars move at a constant velocity in a straight line.
If they move in opposite directions the cars will be 600m apart after 30s.
If they move in the same direction they will be 25m apart after 25s.
Calculate their velocities in m/s.
:
let a = A car's velocity
let b = B car's velocity
:
Write a distance equation for each scenario; dist = time * speec
30(a+b) = 600, relative speed is the sum
25(a-b) = 25, relative speed is the difference
:
Simplify both equations, divide the 1st by 30, the 2nd by 25 and you have
a + b = 20
a - b = 1
---------------adding eliminates b, find a
2a = 21
a = 21/2
a = 10.5 m/sec is A's velocity
Find b
b = 20 - 10.5
b = 9.5 m/sec is B's velocity
:
You can confirm this by putting these values in both original equations